Dear Tutor,
Could you please help me with this problem?
I'm not sure if I selected the right topic.
It says" Solve the system using any algebraic method.
(here is the problem.)
x + 4y + z = 12
y - 3z = -7
z = 3
Thank you so much for helping me.

Truly,
Kat

Answer by slksil001(7)


x + 4y + z = 12.........(A)
y - 3z = -7 ............(B)
z = 3 ..................(C)
subst (C) in B
y= -7+3(3)
=2 ................(D)


subt. (D)&(C) in (A)
x+4(2)+3 = 12
=12-8-3
=1
after you can check that your X ,Y & Z are correct by plugging them back in the equation (A):

1 + 8 + 3 = 12
0 = 0